### v2

@@ Abstract
Applying deep reinforcement learning (DRL) to quadrupedal robots is promising for obstacle negotiation, terrain locomotion, and real-world deployment. Existing methods either train blind agents that sacrifice foresight or adopt cross-modal fusion architectures with limitations; even Transformer-based models still suffer from the quadratic cost of self-attention. To overcome these challenges, we propose LocoMamba, a vision-driven cross-modal DRL framework built on selective state-space models (Mamba). LocoMamba encodes proprioceptive and vision features into compact tokens, which are fused by stacked Mamba layers through near-linear-time selective scanning. This reduces latency and memory usage while preserving long-range dependencies. The policy is trained end-to-end with Proximal Policy Optimization under randomized terrain and obstacle-density curriculum, using a state-centric reward balancing task and safety. Evaluations in diverse static and dynamic obstacle settings show that LocoMamba significantly outperforms a Transformer baseline, achieving $48.9\%$ higher returns, $30.4\%$ longer distance, and $48.9\%$ fewer collisions on trained terrains, while exhibiting stronger generalization to unseen scenarios. Moreover, it converges faster with limited compute, enhancing training efficiency. Overall, LocoMamba improves locomotion performance and reduces training cost, supporting rapid iteration and deployment of quadruped learning algorithms. A repository is hosted at https://github.com/allen-quad-robot/locomamba, to be made available upon acceptance.
@@ 1 Introduction
Quadrupedal robots provide mobility in environments where wheeled platforms are ineffective, such as stairs, rubble, soft or deformable substrates, and cluttered indoor or outdoor settings, enabling applications in inspection, disaster response, agriculture, and planetary exploration [1]. Robust locomotion control is therefore a foundational capability for practical quadrupedal systems, underpinning safe navigation, dependable mission execution, and reliable operation across diverse terrains and disturbance conditions [2].
Current approaches to quadrupedal locomotion control fall into two broad categories. Classical model-based pipelines depend on accurate dynamics and terrain models, explicit contact planning, and extensive parameter tuning, which constrain scalability and complicate deployment [3]. Learning-based methods mitigate this burden and enable end-to-end policies that couple perception with control [4]. Within this paradigm, deep reinforcement learning has emerged as the dominant approach because it optimizes task-driven closed-loop policies through interaction with the environment and produces robust, adaptive behaviors [5].
DRL methods have substantially advanced quadrupedal locomotion, enabling traversal of uneven terrain [6] and operation under challenging conditions such as mud, snow, and running water [7]. However, most studies continue to train blind controllers that rely only on proprioception and achieve stability primarily through large-scale simulation and domain randomization. Despite their robustness, blind agents lack foresight because they receive no exteroceptive input, so they react only after contact and struggle to avoid obstacles proactively or to plan foot placement on irregular ground [8].
On the other hand, cross-modal fusion with visual input has gained traction. Vision complements proprioception by providing look-ahead, enabling early perception of distant obstacles, anticipation of terrain changes before contact, and timely trajectory adjustments that reduce collisions and improve foot placement [9]. Nevertheless, current fusion architectures exhibit notable trade-offs. Recurrent models often suffer from vanishing gradients and limited capacity for long-horizon dependencies [10]. Hierarchical designs complicate optimization and propagate errors across levels [11]. Transformer-based fusion, although expressive, incurs quadratic memory use and computational cost with the number of tokens, which constrains sequence length, spatial resolution, and training efficiency [12].
@@ 2.1 Learning-Based Quadrupedal Locomotion
Research on quadrupedal locomotion has evolved along two principal paradigms: rule-based control and learning-based approaches. Early rule-based controllers relied on template dynamics and heuristic regulation of balance, leg compliance, and foot placement, establishing fundamental principles for dynamic locomotion [13, 14, 15, 16]. Subsequent rule-based systems formalized locomotion as constrained optimal control, including trajectory optimization over centroidal or full-body dynamics and model predictive control with explicit contact constraints and task hierarchies [17, 18, 19, 20]. For example, [21] developed real-time MPC and task-space whole-body control on torque-controlled quadrupeds, demonstrating stable trotting, stair traversal, and robust disturbance rejection. [22] refined MPC cost design and contact modeling to improve tracking and compliance under high-frequency control budgets. [23] studied trajectory optimization and contact-aware planning for foothold selection and force allocation, improving reliability on structured terrain. However, these approaches require accurate models of the environment and system dynamics and substantial manual tuning, thereby limiting their applicability in complex and variable settings.
On the other hand, model-free RL can learn general policies for challenging conditions [24, 25, 26, 4]. For example, [27] trained torque-level policies with dynamics randomization and privileged learning to achieve fast, robust trotting on uneven terrain. [28] used large-scale domain randomization and dynamics perturbations to transfer simulation-trained policies to real quadrupeds with strong disturbance rejection. [29] introduced rapid motor adaptation that augments a proprioceptive policy with an online adaptation module, enabling quick recovery under payload shifts and terrain changes. However, most approaches are proprioception-only and thus blind to exteroceptive cues, which limits foresight for obstacle negotiation and precise foothold planning. In this work, both the visual and proprioceptive inputs are fused to obtain a richer state representation, enabling the policy to anticipate terrain changes and plan trajectories online while maintaining continuous motion.
@@ 2.2 Vision-Driven RL for Quadrupedal Locomotion
To extend RL beyond state-only inputs, a growing body of work leveraged visual observations for locomotion control [9]. For example, Yu et al. [30] trained a vision-based controller that incorporated exteroceptive input directly into the RL loop and achieved traversal of uneven terrain and complex obstacles. Duan et al. [31] learned a controller in simulation with heightmap observations and then trained a depth-to-heightmap predictor from depth and state histories, enabling vision-guided locomotion with transfer to hardware on challenging terrains. Complementing these efforts, Fahmi et al. [32] proposed the ViTAL framework, which used vision for terrain-aware planning by decomposing decision making into foothold selection and pose adaptation, and improved safety and reliability on irregular surfaces. Hierarchical formulations were also explored; for example, Jain et al. [11] employed hierarchical RL in which high-level visual policies guided low-level motor control with sensor inputs.
Despite these advances, common fusion strategies involved trade-offs. Simple concatenation underused spatial structure, and hierarchical decompositions increased optimization complexity and risked error propagation across levels [33]. In this work, complex hierarchical designs are avoided and a simple end-to-end schema is adopted that fuses proprioceptive and depth inputs within a single policy trained with PPO, enabling visual look-ahead while maintaining a streamlined learning pipeline.
@@ 2.3 State-Space Models and Cross-Modal RL
Cross-modal RL has attracted significant attention. Typical fusion backbones include recurrent neural networks (RNNs) and Transformers. For example, [34] designed an egocentric visuomotor RL framework that fused egocentric depth and proprioception with long short-term memory (LSTM) units, enabling a quadruped to negotiate obstacles and traverse cluttered scenes in the real world. [35] proposed a gated recurrent unit (GRU)–based world-model formulation that fuses depth images and proprioceptive readings into a recurrent latent state for visual legged locomotion, improving control under partial observability. [12] introduced a cross-modal Transformer policy that ingests depth tokens and proprioceptive features end-to-end, achieving stronger terrain anticipation and sim-to-real transfer than blind baselines; subsequent work scaled such tokenized proprioception-vision Transformer policies to heterogeneous robot datasets to improve generalization. However, RNNs often suffer from vanishing gradients and limited long-horizon capacity, which complicates optimization in extended sequences. Transformer models offer strong expressivity but incur quadratic memory and computation with the number of tokens, which constrains sequence length, spatial resolution, and training efficiency.
SSMs [36] offer an alternative by updating a compact recurrent state per token and scaling near-linearly with sequence length. Building on the S4 model [37], which demonstrated efficient long-range dependency modeling on long-sequence benchmarks, [38] introduced Mamba , a selective SSM whose input-dependent parameters and hardware-aware parallel scan yield linear-time sequence modeling with favorable throughput and memory characteristics across modalities.
Mamba-style SSMs have also been explored for multimodal fusion. AlignMamba augments a multimodal Mamba backbone with token-level optimal-transport alignment and a global distributional alignment objective to improve cross-modal consistency [39]. COMO (Cross-Mamba Interaction with Offset-Guided Fusion) addresses sensor misalignment for multimodal object detection while retaining the efficiency benefits of selective scanning [40]. AV-Mamba applies selective SSMs to audio-visual question answering and reports improvements over Transformer baselines [41]. FusionMamba [42] and DepMamba [43] further demonstrate SSM-based gains in multimodal image fusion and audio-visual affect analysis. To the best of our knowledge, this is the first work to use cross-modal Mamba for quadrupedal locomotion.
@@ 3 Methodology
Fig. 1 presents the overall architecture of LocoMamba. The pipeline comprises three components trained end to end. First, proprioceptive states and depth images are encoded into a compact tokenized latent space: an MLP maps the proprioceptive vector to a state token, and a lightweight CNN patchifies the depth image into spatial tokens. Second, the concatenated token stream is fused by stacked Mamba SSM layers via selective state-space scanning, which updates a compact recurrent state per token, yields near linear time and memory scaling, accommodates variable token counts and resolutions, and preserves long-horizon temporal context. Third, policy and value heads consume the fused features and are optimized with PPO [44] under a compact state-centric reward that balances task-aligned progress, energy efficiency, and safety.
@@ 3.1 Proprioception–Depth Input Encoding
The depth image is processed by a compact CNN [45] to produce a sequence of spatial tokens:
@@ 3.2 State Space Modelling with Mamba
The cross-modal token sequence $\hat{U}_{t}$ is fused with a stack of Mamba SSM layers [38] (Fig. 2). Each layer scans the token stream with input-dependent state updates and carries a compact recurrent state across tokens and time steps, enabling efficient long-horizon modeling.
@@
where $\bar{A}_{t,k},\bar{B}_{t,k},\bar{C}_{t,k},\bar{D}_{t,k}$ are input-gated, token-dependent parameters produced by lightweight affine transforms of $u_{t,k}$. The scan proceeds over $k{=}1\ldots(1{+}N)$, and the initial state at each time step is carried from the previous step, $x_{t,1}\leftarrow x_{t-1,\,1+N}$. The output tokens of the layer are $Y_{t}=\{y_{t,k}\}_{k=1}^{1+N}\in\mathbb{R}^{(1+N)\times d}$. The token width $h$ and the hidden state dimension of the SSM $d$ are set to equal by a learned projection when needed.
@@
with $H^{(0)}_{t}\!=\!\hat{U}_{t}$ and $Y^{(\ell)}_{t}$ computed by (9)–(10) using inputs $H^{(\ell)}_{t}$. The selective SSM update yields near-linear time and memory in the token count (1+N)(1{+}N), in contrast to the quadratic cost of global self-attention [33], and naturally supports streaming over time by reusing the carried state $x_{t,1}$.
@@ 3.3 Policy Optimization with PPO
The locomotion can be modeled as a discounted MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,\gamma)$ with discount $\gamma\in(0,1)$[46]. At time $t$, the agent receives observation $o_{t}$ from environment and obtains a fused feature $h_{t}=f_{\text{fuse}}(o_{t})\in\mathbb{R}^{d_{h}}$ from the Mamba backbone (Sections 3.2). The objective is to maximize the expected discounted return
@@
The policy is optimized with PPO [44] using the clipped surrogate. With likelihood ratio
@@ 4.1 Environment Setup
All experiments are conducted on a laptop equipped with an Intel Core Ultra 7 155H CPU (22 cores, 1.4 GHz base clock) and an NVIDIA GeForce RTX 500 Ada GPU (4 GB, CUDA-12.9). The operating system is Ubuntu 22.04. The physics simulation is run with PyBullet [47] and all models are implemented in Python 3.8 with PyTorch 2.4.1.
@@ 4.2 RL MDP Details
The action is a 12-dimensional vector that specifies desired joint position targets for the 12 actuated joints. A proportional–derivative (PD) controller [48] converts these targets into joint torques on the agent.
@@
where $\alpha_{\text{fwd}}{=}1$, $\alpha_{\text{energy}}{=}0.005$, $\alpha_{\text{alive}}{=}0.1$, and $K_{t}$ is the number of collectible spheres obtained at step $t$ (if applicable). The forward term encourages task-aligned motion,

### v3

@@ Abstract
Applying deep reinforcement learning (DRL) to quadrupedal robots is promising for obstacle negotiation, terrain locomotion, and real-world deployment. Existing methods either train blind agents that sacrifice foresight or adopt cross-modal fusion architectures with limitations; even Transformer-based models still suffer from the quadratic cost of self-attention. To overcome these challenges, we propose LocoMamba, a vision-driven cross-modal DRL framework built on selective state-space models (Mamba). LocoMamba encodes proprioceptive and vision features into compact tokens, which are fused by stacked Mamba layers through near-linear-time selective scanning. This reduces latency and memory usage while preserving long-range dependencies. The policy is trained end-to-end with Proximal Policy Optimization under randomized terrain and obstacle-density curriculum, using a state-centric reward balancing task and safety. Evaluations in diverse static and dynamic obstacle settings show that LocoMamba significantly outperforms a Transformer baseline, achieving $48.9\%$ higher returns, $30.4\%$ longer distance, and $48.9\%$ fewer collisions on trained terrains, while exhibiting stronger generalization to unseen scenarios. Moreover, it converges faster with limited compute, enhancing training efficiency. Overall, LocoMamba improves locomotion performance and reduces training cost, supporting rapid iteration and deployment of quadruped learning algorithms. A repository is hosted at https://github.com/allen-quad-robot/locomamba.
@@ 1 Introduction
Quadrupedal robots provide mobility in environments where wheeled platforms are ineffective, such as stairs, rubble, soft or deformable substrates, and cluttered indoor or outdoor settings, enabling applications in inspection, disaster response, agriculture, and planetary exploration [fan2024review]. Robust locomotion control is therefore a foundational capability for practical quadrupedal systems, underpinning safe navigation, dependable mission execution, and reliable operation across diverse terrains and disturbance conditions [carpentier2021recent].
Current approaches to quadrupedal locomotion control fall into two broad categories. Classical model-based pipelines depend on accurate dynamics and terrain models, explicit contact planning, and extensive parameter tuning, which constrain scalability and complicate deployment [xin2021robust]. Learning-based methods mitigate this burden and enable end-to-end policies that couple perception with control [ha2025learning]. Within this paradigm, deep reinforcement learning has emerged as the dominant approach because it optimizes task-driven closed-loop policies through interaction with the environment and produces robust, adaptive behaviors [zhang2022deepreinforcementlearningforreal].
DRL methods have substantially advanced quadrupedal locomotion, enabling traversal of uneven terrain [xie2021dynamics] and operation under challenging conditions such as mud, snow, and running water [lee2020learning]. However, most studies continue to train blind controllers that rely only on proprioception and achieve stability primarily through large-scale simulation and domain randomization. Despite their robustness, blind agents lack foresight because they receive no exteroceptive input, so they react only after contact and struggle to avoid obstacles proactively or to plan foot placement on irregular ground [xie2022glide].
On the other hand, cross-modal fusion with visual input has gained traction. Vision complements proprioception by providing look-ahead, enabling early perception of distant obstacles, anticipation of terrain changes before contact, and timely trajectory adjustments that reduce collisions and improve foot placement [han2025multimodal]. Nevertheless, current fusion architectures exhibit notable trade-offs. Recurrent models often suffer from vanishing gradients and limited capacity for long-horizon dependencies [li2024rapid]. Hierarchical designs complicate optimization and propagate errors across levels [jain2019hierarchical]. Transformer-based fusion, although expressive, incurs quadratic memory use and computational cost with the number of tokens, which constrains sequence length, spatial resolution, and training efficiency [yang2021learning].
@@ 2.1 Learning-Based Quadrupedal Locomotion
Research on quadrupedal locomotion has evolved along two principal paradigms: model-based optimal control and learning-based approaches. Early model-based optimal control relied on template dynamics and heuristic regulation of balance, leg compliance, and foot placement, establishing fundamental principles for dynamic locomotion [miura1984dynamic, liu2014planning, habu2018simple, bledt2018cheetah]. Subsequent model-based optimal control systems formalized locomotion as constrained optimal control, including trajectory optimization over centroidal or full-body dynamics and model predictive control with explicit contact constraints and task hierarchies [grandia2019feedback, yao2021hierarchy, amatucci2024accelerating, elobaid2025adaptive]. For example, [di2018dynamic] developed real-time MPC and task-space whole-body control on torque-controlled quadrupeds, demonstrating stable trotting, stair traversal, and robust disturbance rejection. [ding2019real] refined MPC cost design and contact modeling to improve tracking and compliance under high-frequency control budgets. [carius2019trajectory] studied trajectory optimization and contact-aware planning for foothold selection and force allocation, improving reliability on structured terrain. However, these approaches require accurate models of the environment and system dynamics and substantial manual tuning, thereby limiting their applicability in complex and variable settings.
On the other hand, model-free RL can learn general policies for challenging conditions [li2021reinforcement, margolis2024rapid, bussola2025guided, ha2025learning]. For example, [hwangbo2019learning] trained torque-level policies with dynamics randomization and privileged learning to achieve fast, robust trotting on uneven terrain. [tan2018sim] used large-scale domain randomization and dynamics perturbations to transfer simulation-trained policies to real quadrupeds with strong disturbance rejection. [kumar2021rma] introduced rapid motor adaptation that augments a proprioceptive policy with an online adaptation module, enabling quick recovery under payload shifts and terrain changes. However, most approaches are proprioception-only and thus blind to exteroceptive cues, which limits foresight for obstacle negotiation and precise foothold planning. In this work, both the visual and proprioceptive inputs are fused to obtain a richer state representation, enabling the policy to anticipate terrain changes and plan trajectories online while maintaining continuous motion.
@@ 2.2 Vision-Driven RL for Quadrupedal Locomotion
To extend RL beyond state-only inputs, a growing body of work leveraged visual observations for locomotion control [han2025multimodal]. For example, yu2021visual trained a vision-based controller that incorporated exteroceptive input directly into the RL loop and achieved traversal of uneven terrain and complex obstacles. duan2024learning learned a controller in simulation with heightmap observations and then trained a depth-to-heightmap predictor from depth and state histories, enabling vision-guided locomotion with transfer to hardware on challenging terrains. Complementing these efforts, fahmi2022vital proposed the ViTAL framework, which used vision for terrain-aware planning by decomposing decision making into foothold selection and pose adaptation, and improved safety and reliability on irregular surfaces. Hierarchical formulations were also explored; for example, jain2019hierarchical employed hierarchical RL in which high-level visual policies guided low-level motor control with sensor inputs.
Despite these advances, common fusion strategies involved trade-offs. Simple concatenation underused spatial structure, and hierarchical decompositions increased optimization complexity and risked error propagation across levels [singh2022reinforcement]. In this work, complex hierarchical designs are avoided and a simple end-to-end schema is adopted that fuses proprioceptive and depth inputs within a single policy trained with PPO, enabling visual look-ahead while maintaining a streamlined learning pipeline.
@@ 2.3 State-Space Models and Cross-Modal RL
Cross-modal RL has attracted significant attention. Typical fusion backbones include recurrent neural networks (RNNs) and Transformers. For example, [xiao2024egocentric] designed an egocentric visuomotor RL framework that fused egocentric depth and proprioception with long short-term memory (LSTM) units, enabling a quadruped to negotiate obstacles and traverse cluttered scenes in the real world. [lai2024world] proposed a gated recurrent unit (GRU)–based world-model formulation that fuses depth images and proprioceptive readings into a recurrent latent state for visual legged locomotion, improving control under partial observability. [yang2021learning] introduced a cross-modal Transformer policy that ingests depth tokens and proprioceptive features end-to-end, achieving stronger terrain anticipation and sim-to-real transfer than blind baselines; subsequent work scaled such tokenized proprioception-vision Transformer policies to heterogeneous robot datasets to improve generalization. However, RNNs often suffer from vanishing gradients and limited long-horizon capacity, which complicates optimization in extended sequences. Transformer models offer strong expressivity but incur quadratic memory and computation with the number of tokens, which constrains sequence length, spatial resolution, and training efficiency.
SSMs [kalman1960new] offer an alternative by updating a compact recurrent state per token and scaling near-linearly with sequence length. Building on the S4 model [gu2021efficiently], which demonstrated efficient long-range dependency modeling on long-sequence benchmarks, [gu2023mamba] introduced Mamba , a selective SSM whose input-dependent parameters and hardware-aware parallel scan yield linear-time sequence modeling with favorable throughput and memory characteristics across modalities.
Mamba-style SSMs have also been explored for multimodal fusion. AlignMamba augments a multimodal Mamba backbone with token-level optimal-transport alignment and a global distributional alignment objective to improve cross-modal consistency [li2025alignmamba]. COMO (Cross-Mamba Interaction with Offset-Guided Fusion) addresses sensor misalignment for multimodal object detection while retaining the efficiency benefits of selective scanning [liu2026cross]. AV-Mamba applies selective SSMs to audio-visual question answering and reports improvements over Transformer baselines [huang2024av]. FusionMamba [xie2024fusionmamba] and DepMamba [ye2025depmamba] further demonstrate SSM-based gains in multimodal image fusion and audio-visual affect analysis. Unlike prior cross-modal Mamba variants that focus on perception-level multimodal alignment, our work investigates its benefits for robotic locomotion control, where real-time decision latency, causal sensor fusion, and long-horizon dependency are critical. This setting highlights Mamba’s suitability for efficient and temporally consistent proprioception–vision fusion within an end-to-end RL framework. To the best of our knowledge, this is the first work to use cross-modal Mamba for quadrupedal locomotion.
@@ 3 Methodology
Fig. 1 presents the overall architecture of LocoMamba. The pipeline comprises three components trained end to end. First, proprioceptive states and depth images are encoded into a compact tokenized latent space: an MLP maps the proprioceptive vector to a state token, and a lightweight CNN patchifies the depth image into spatial tokens. Second, the concatenated token stream is fused by stacked Mamba SSM layers via selective state-space scanning, which updates a compact recurrent state per token, yields near linear time and memory scaling, accommodates variable token counts and resolutions, and preserves long-horizon temporal context. Third, policy and value heads consume the fused features and are optimized with PPO [schulman2017proximal] under a compact state-centric reward that balances task-aligned progress, energy efficiency, and safety.
@@ 3.1 Proprioception–Depth Input Encoding
The depth image is processed by a compact CNN [li2021survey] to produce a sequence of spatial tokens:
@@ 3.2 State Space Modelling with Mamba
The cross-modal token sequence $\hat{U}_{t}$ is fused with a stack of Mamba SSM layers [gu2023mamba] (Fig. 2). Each layer scans the token stream with input-dependent state updates and carries a compact recurrent state across tokens and time steps, enabling efficient long-horizon modeling.
@@
where $\bar{A}_{t,k},\bar{B}_{t,k},\bar{C}_{t,k},\bar{D}_{t,k}$ are input-gated, token-dependent parameters produced by lightweight affine transforms of $u_{t,k}$. The scan proceeds over $k{=}1\ldots(1{+}N)$, and the initial state at each time step is carried from the previous step, $x_{t,1}\leftarrow x_{t-1,\,1+N}$. The output tokens of the layer are $Y_{t}=\{y_{t,k}\}_{k=1}^{1+N}\in\mathbb{R}^{(1+N)\times d}$. The token width $h$ and the hidden state dimension of the SSM $d$ are set to equal by a learned projection when needed. In this work, a two-layer stacked Mamba backbone is adopted as a balanced configuration that captures both short- and mid-term temporal dependencies while maintaining training stability and computational efficiency.
@@
with $H^{(0)}_{t}\!=\!\hat{U}_{t}$ and $Y^{(\ell)}_{t}$ computed by (9)–(10) using inputs $H^{(\ell)}_{t}$. The selective SSM update yields near-linear time and memory in the token count (1+N)(1{+}N), in contrast to the quadratic cost of global self-attention [singh2022reinforcement], and naturally supports streaming over time by reusing the carried state $x_{t,1}$.
@@ 3.3 Policy Optimization with PPO
The locomotion can be modeled as a discounted MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,\gamma)$ with discount $\gamma\in(0,1)$[richard1957mark]. At time $t$, the agent receives observation $o_{t}$ from environment and obtains a fused feature $h_{t}=f_{\text{fuse}}(o_{t})\in\mathbb{R}^{d_{h}}$ from the Mamba backbone (Sections 3.2). The objective is to maximize the expected discounted return
@@
The policy is optimized with PPO [schulman2017proximal] using the clipped surrogate. With likelihood ratio
@@ 4.1 Environment Setup
All experiments are conducted on a laptop equipped with an Intel Core Ultra 7 155H CPU (22 cores, 1.4 GHz base clock) and an NVIDIA GeForce RTX 500 Ada GPU (4 GB, CUDA-12.9). The operating system is Ubuntu 22.04. The physics simulation is run with PyBullet [coumans2021pybullet] and all models are implemented in Python 3.8 with PyTorch 2.4.1.
@@ 4.2 RL MDP Details
The action is a 12-dimensional vector that specifies desired joint position targets for the 12 actuated joints. A proportional–derivative (PD) controller [tan2009computation] converts these targets into joint torques on the agent.
@@
where $\alpha_{\text{fwd}}{=}1$, $\alpha_{\text{energy}}{=}0.005$, $\alpha_{\text{alive}}{=}0.1$.
The forward term encourages task-aligned motion,